ON THE CHARACTERISTIC PROPERTIES OF EXPONENTIAL DISTRIBUTION

ISMIHAN G. BAIRAMOV

Department of Statistics, Ankara University, 06100, Tandogan, Ankara, Turkey

(Received April 6, 1998; revised April 16, 1999)

Abstract. New characterizations for the exponential distribution are given in terms of record values and the probabilities of finite sums of independent and identically distributed nonnegative random variables provided that the underlying distribution is either new better than used or new worse than used.

Key words and phrases: Characterization, exponential distribution, order statistics, record value, new better than used.